# Truthfulness of Calibration Measures

### v1

@@ Abstract
We initiate the study of the truthfulness of calibration measures in sequential prediction. A calibration measure is said to be truthful if the forecaster (approximately) minimizes the expected penalty by predicting the conditional expectation of the next outcome, given the prior distribution of outcomes. Truthfulness is an important property of calibration measures, ensuring that the forecaster is not incentivized to exploit the system with deliberate poor forecasts. This makes it an essential desideratum for calibration measures, alongside typical requirements, such as soundness and completeness.
@@ 1 Introduction
While truthfulness of calibration measures has not been systematically investigated to date, evidence of the lack of truthfulness of some calibration measures has emerged in recent literature (e.g., [FH21, QV21]). More recently, [QZ24] highlighted a large gap in the truthfulness of a recently proposed calibration measure (called the distance from calibration [BGHN23]) by showing that in a simple setup of predicting i.i.d. outcomes, the truthful forecaster incurs a distance of $\Omega(\sqrt{T})$ from calibration but there is a forecasting algorithm that achieves $\operatorname{polylog}(T)$ distance from calibration. We call this a $\operatorname{polylog}(T)$-$\Omega(\sqrt{T})$ truthfulness gap. On the other hand, we say that a calibration measure is $(\alpha,\beta)$-truthful if predicting the next outcome according to its conditional distribution incurs a measure that is no more than $\alpha\mathsf{OPT}+\beta$, where $\mathsf{OPT}$ is the minimum value of the calibration measure achievable by any forecaster. Faced with evidence that some calibration measures suffer from large truthfulness gaps, we will systematically examine the truthfulness (or a gap thereof) of a wide range of calibration measures.
@@
We do this using an embarrassingly simple adjustment to an existing notion of calibration measure: we subsample a subset of the time steps and evaluate the smooth calibration error [KF08] on this sampled set only. We call this the Subsampled Smooth Calibration Error (SSCE) and formally define it in Section 2. Our main result is that SSCE is $(O(1),0)$-truthful.
@@ 1.1 Related Work
There is a large body of work on calibration, a notion that dates back to the 1950s [Bri50, Daw82, Daw85] and has been applied to game theory [FV97], machine learning [GPSW17], and algorithmic fairness [KMR17, PRW+17, HJKRR18]. We will restrict our discussion to sequential calibration and the systematic study of calibration measures, which are the closest to this work.
@@
Foster and Vohra [FV98] first proved that one can achieve asymptotic calibration on arbitrary and adaptive outcomes. Formally, they gave a forecasting algorithm with an $O(T^{2/3})$ ECE in expectation, when predicting $T$ binary outcomes chosen by an adaptive adversary. Subsequent work gave alternative and simpler proofs of the result [FL99, Fos99, Har22], extended the result to other calibration measures [KF08, FH18, FH21, QZ24], and proved lower bounds on the optimal ECE [QV21, DDF+24]. Most closely related to our approach is the work of [FRST11], who studied a stronger notion that requires calibration on a family of checking rules, where each checking rule specifies a subset of the time horizon. Despite the apparent similarity, their notion is qualitatively different from the SSCE, since we take an expectation over the subsampled horizon, whereas they take the maximum. In particular, no forecaster can be calibrated in their definition if the checking rule family contains all subsets of $[T]$, since there always exists a checking rule that strongly correlates with the outcomes.
@@ 2 Preliminaries
The quality of the forecaster’s predictions in the setting above is quantified by calibration measures. Formally, a calibration measure $\mathsf{CM}$ is a family of functions $\{\mathsf{CM}_{T}:T\in\mathbb{N}\}$, where each $\mathsf{CM}_{T}$ maps $\{0,1\}^{T}\times[0,1]^{T}$ to $[0,T]$. We will frequently omit the subscript $T$, since it is usually clear from the context. With respect to calibration measure $\mathsf{CM}$, the expected penalty incurred by forecaster $\mathcal{A}$ on distribution $\mathcal{D}$ is defined as $\mathsf{err}_{\mathsf{CM}}(\mathcal{D},\mathcal{A})\coloneqq\operatorname*{\mathbb{E}}_{(x,p)\sim(\mathcal{D},\mathcal{A})}\left[\mathsf{CM}(x,p)\right]$, where $(x,p)\sim(\mathcal{D},\mathcal{A})$ denotes sampling a sequence $x$ and predictions $p$ from the joint distribution induced by $\mathcal{D}$ and $\mathcal{A}$.
@@
$\mathsf{SSCE}(x,p)\coloneqq\operatorname*{\mathbb{E}}_{S\sim\mathsf{Unif}(2^{[T]})}\left[\mathsf{smCE}(x|_{S},p|_{S})\right]=\operatorname*{\mathbb{E}}_{y\sim\mathsf{Unif}(\{0,1\}^{T})}\left[\sup_{f\in\mathcal{F}}\sum_{t=1}^{T}y_{t}\cdot f(p_{t})\cdot(x_{t}-p_{t})\right].$
@@
With respect to distribution $\mathcal{D}\in\Delta(\{0,1\}^{T})$, the truthful forecaster is defined as $\mathcal{A}^{\mathrm{truthful}}(\mathcal{D})(b_{1},b_{2},\ldots,b_{t-1})\coloneqq\Pr_{x\sim\mathcal{D}}\left[x_{t}=1\ \Big{|}\ x_{1:(t-1)}=b_{1:(t-1)}\right]$.
@@
The optimal error on distribution $\mathcal{D}\in\Delta(\{0,1\}^{T})$ with respect to calibration measure $\mathsf{CM}$ is defined as $\mathsf{OPT}_{\mathsf{CM}}(\mathcal{D})\coloneqq\inf_{\mathcal{A}}\mathsf{err}_{\mathsf{CM}}(\mathcal{D},\mathcal{A})$, where $\mathcal{A}$ ranges over all deterministic forecasters.
@@ 3 Technical Overview
As we elaborate in Section 5, the crux of Step (2) is to control the expected deviation of a martingale $(M_{t})_{0\leq t\leq T}$ with respect to filtration $(\mathbb{F}_{t})$ by the its realized variance $\mathrm{Var}_{t}\coloneqq\sum_{s=1}^{t}\operatorname*{\mathrm{Var}}\left[M_{s}|\mathbb{F}_{s-1}\right]$, which is highly non-trivial as the two processes $(M_{t})$ and $(\mathrm{Var}_{t})$ are correlated. In more detail, the filtration $(\mathbb{F}_{t})$ corresponds to the randomness in $x\sim\mathcal{D}$, while $(M_{t})$ tracks the biases in the predictions (on a subset of the time horizon) tested by a Lipschitz function. We note that such a bound would easily follow from “off-the-shelf” concentration inequalities for martingales (e.g., Freedman’s inequality [Fre75]), if the total realized variance $\mathrm{Var}_{T}$ were uniformly bounded. However, in general, $\mathrm{Var}_{T}$ may vary drastically, and directly applying these concentration inequalities would introduce an extra super-constant factor. Our workaround is a “doubling trick” that divides the time horizon into epochs, the realized variances in which grow exponentially. We then apply Freedman’s inequality to each epoch separately. In Section 5, we formulate a toy random walk problem that highlights this challenge and demonstrates our solution to it, which is of independent interest.
Similarly, as we show in Section 6, the crux of Step (3) is to establish another martingale inequality. We first show that for fixed $x$ and $p$, we have $\mathsf{SSCE}(x,p)=\Omega(\sqrt{N_{T}})$, where $N_{t}\coloneqq\sum_{s=1}^{t}\mathbbm{1}\left[|x_{s}-p_{s}|\geq 1/2\right]$. Furthermore, over the randomness in $x\sim\mathcal{D}$, the realized variance process $(\mathrm{Var}_{t})$ defined above is shown to lower bound $(N_{t})$, i.e., $(N_{t}-\mathrm{Var}_{t})$ is a sub-martingale. However, the desired result requires the lower bound $\operatorname*{\mathbb{E}}\left[\sqrt{N_{T}}\right]\geq\Omega(1)\cdot\operatorname*{\mathbb{E}}\left[\sqrt{\mathrm{Var}_{T}}\right]$, which does not follow from $\operatorname*{\mathbb{E}}\left[N_{T}-\mathrm{Var}_{T}\right]\geq 0$ in general. This challenge necessitates a more careful analysis tailored to the specific properties of the processes $(N_{t})$ and $(\mathrm{Var}_{t})$.
@@ 4 Warmup: The Product Distribution Case
For distribution $\mathcal{D}=\prod_{t=1}^{T}\mathsf{Bernoulli}(p^{\star}_{t})$, take $\sigma^{2}\coloneqq\operatorname*{\mathrm{Var}}_{x\sim\mathcal{D}}\left[\sum_{t=1}^{T}x_{t}\right]=\sum_{t=1}^{T}p^{\star}_{t}(1-p^{\star}_{t})$ as a complexity measure of the distribution of outcomes. We will show that $\mathsf{err}_{\mathsf{SSCE}}(\mathcal{D},\mathcal{A}^{\mathrm{truthful}}(\mathcal{D}))=O(\sigma+\log T)$ and $\mathsf{OPT}_{\mathsf{SSCE}}(\mathcal{D})=\Omega(\sigma)-O(1)$.
@@ 4.1 Upper Bound the SSCE of the Truthful Forecaster
where $\mathcal{G}_{\delta}\coloneqq\{f_{\delta}-f_{\delta/2}:f_{\delta}\in\mathcal{F}_{\delta},f_{\delta/2}\in\mathcal{F}_{\delta/2},\|f_{\delta}-f_{\delta/2}\|_{\infty}\leq 3\delta/2\}$.
@@ 5 Upper Bound the SSCE of the Truthful Forecaster
To extend the proof strategy sketched in Section 4 to non-product distributions, the first challenge is to define an appropriate complexity measure of a general distribution $\mathcal{D}$. Consider the stochastic process $(\mathrm{Var}_{t})_{0\leq t\leq T}$ defined as $\mathrm{Var}_{t}\coloneqq\sum_{s=1}^{t}p^{\star}_{s}(1-p^{\star}_{s})$, where
$p^{\star}_{t}\coloneqq\operatorname*{\mathbb{E}}_{x^{\prime}\sim\mathcal{D}}\left[x^{\prime}_{t}\Big{|}x^{\prime}_{1:(t-1)}=x_{1:(t-1)}\right]$
@@
To control the expectation of each $W_{k}$, we note that the set $\mathcal{G}_{k}\coloneqq\left\{\pi_{2^{-k}}(f)-\pi_{2^{1-k}}(f):f\in\mathcal{F}\right\}$ is of size at most $|\mathcal{F}_{2^{-k}}|\cdot|\mathcal{F}_{2^{1-k}}|$. Furthermore, every function $g\in\mathcal{G}_{k}$ satisfies
@@
Given a function $f:[0,1]\to[-1,1]$ and $y\in\{0,1\}^{T}$, consider the martingale $M_{t}(f,y)\coloneqq\sum_{s=1}^{t}y_{s}\cdot f(p^{\star}_{s})\cdot(x_{s}-p^{\star}_{s})$ where $x\sim\mathcal{D}$. Then, for any finite family $\mathcal{G}$ of functions from $[0,1]$ to $[-1,1]$ and any $y\in\{0,1\}^{T}$, we have
@@
$\mathrm{Var}_{t}(\mathcal{I})\coloneqq\sum_{s=1}^{t}p^{\star}_{s}(1-p^{\star}_{s})\cdot\mathbbm{1}\left[p^{\star}_{s}\in\mathcal{I}\right],$
where $x\sim\mathcal{D}$ and $p^{\star}_{t}\coloneqq\Pr_{x^{\prime}\sim\mathcal{D}}\left[x^{\prime}_{t}=1|x^{\prime}_{1:(t-1)}=x_{1:(t-1)}\right]$. We define the epochs with respect to $\mathcal{I}$ as the sequence $\tau_{0},\tau_{1},\dots\in\mathbb{N}$ where $\tau_{0}=0$ and, for each $k\in[\left\lceil\log_{2}(T)\right\rceil+2]$,
@@ 5.2 Towards a Stronger Upper Bound
Given a function $f:[0,1]\to[-1,1]$, $y\in\{0,1\}^{T}$, and set $\mathcal{I}\subseteq[0,1]$, consider the martingale $M_{t}(f,y,\mathcal{I})\coloneqq\sum_{s=1}^{t}y_{t}\cdot f(p^{\star}_{s})\cdot(x_{s}-p^{\star}_{s})\cdot\mathbbm{1}\left[p^{\star}_{s}\in\mathcal{I}\right]$, where $x\sim\mathcal{D}$, and $p_{t}^{\star}=\Pr_{x^{\prime}\sim\mathcal{D}}\left[x^{\prime}_{t}=1|x^{\prime}_{1:(t-1)}=x_{1:(t-1)}\right]$. Then, for any finite family $\mathcal{G}$ of functions from $[0,1]$ to $[-1,1]$, any $y\in\{0,1\}^{T}$, and any $\mathcal{I}\subseteq[0,1]$, we have

### v2

@@ Abstract
We initiate the study of truthfulness of calibration measures in sequential prediction. A calibration measure is said to be truthful if the forecaster (approximately) minimizes the expected penalty by predicting the conditional expectation of the next outcome, given the prior distribution of outcomes. Truthfulness is an important property of calibration measures, ensuring that the forecaster is not incentivized to exploit the system with deliberate poor forecasts. This makes it an essential desideratum for calibration measures, alongside typical requirements, such as soundness and completeness.
@@ 1 Introduction
While truthfulness of calibration measures has not been systematically investigated to date, evidence of the lack of truthfulness of some calibration measures has emerged in recent literature. For example, [FH21, QV21] noted that a forecaster can lower their ECE by predicting according to the past. This observation was applied in the algorithm of [FH21] and motivated the “sidestepping” technique in the lower bound proof of [QV21]. More recently, [QZ24] highlighted a large gap in the truthfulness of a recently proposed calibration measure (called the distance from calibration [BGHN23]) by showing that in a simple setup of predicting i.i.d. outcomes, the truthful forecaster incurs a distance of $\Omega(\sqrt{T})$ from calibration but there is a forecasting algorithm that achieves $\operatorname{polylog}(T)$ distance from calibration. We call this a $\operatorname{polylog}(T)$-$\Omega(\sqrt{T})$ truthfulness gap. On the other hand, we say that a calibration measure is $(\alpha,\beta)$-truthful if predicting the next outcome according to its conditional distribution incurs a measure that is no more than $\alpha\mathsf{OPT}+\beta$, where $\mathsf{OPT}$ is the minimum value of the calibration measure achievable by any forecaster. Faced with evidence that some calibration measures suffer from large truthfulness gaps, we will systematically examine the truthfulness (or a gap thereof) of a wide range of calibration measures.
@@
We do this using a simple adjustment to an existing notion of calibration measure: we subsample a subset of the time steps and evaluate the smooth calibration error [KF08] on this sampled set only. We call this the Subsampled Smooth Calibration Error (SSCE) and formally define it in Section 2. Our main result is that SSCE is $(O(1),0)$-truthful.
@@ 1.1 Related Work
There is a large body of work on calibration, a notion that dates back to the 1950s [Bri50, Daw82, Daw85] and has been applied to game theory [FV97, HPY23], machine learning [GPSW17], and algorithmic fairness [KMR17, PRW+17, HJKRR18, HJZ23]. We will restrict our discussion to sequential calibration and the systematic study of calibration measures, which are the closest to this work.
@@
Foster and Vohra [FV98] first proved that one can achieve asymptotic calibration on arbitrary and adaptive outcomes. Formally, they gave a forecasting algorithm with an $O(T^{2/3})$ ECE in expectation, when predicting $T$ binary outcomes chosen by an adaptive adversary. Subsequent work gave alternative and simpler proofs of the result [FL99, Fos99, Har22], extended the result to other calibration measures [KF08, FH18, FH21, QZ24], and proved lower bounds on the optimal ECE [QV21]. Most closely related to our approach is the work of [FRST11], who studied a stronger notion that requires calibration on a family of checking rules, where each checking rule specifies a subset of the time horizon. Despite the apparent similarity, their notion is qualitatively different from the SSCE, since we take an expectation over the subsampled horizon, whereas they take the maximum. In particular, no forecaster can be calibrated in their definition if the checking rule family contains all subsets of $[T]$, since there always exists a checking rule that strongly correlates with the outcomes.
@@ 2 Preliminaries
The quality of the forecaster’s predictions in the setting above is quantified by calibration measures. Formally, a calibration measure $\mathsf{CM}$ is a family of functions $\{\mathsf{CM}_{T}:T\in\mathbb{N}\}$, where each $\mathsf{CM}_{T}$ maps $\{0,1\}^{T}\times[0,1]^{T}$ to $[0,T]$. We will frequently omit the subscript $T$, since it is usually clear from the context. With respect to calibration measure $\mathsf{CM}$, the expected penalty incurred by forecaster $\mathcal{A}$ on distribution $\mathcal{D}$ is defined as $\mathsf{err}_{\mathsf{CM}}(\mathcal{D},\mathcal{A})\coloneq\operatorname*{\mathbb{E}}_{(x,p)\sim(\mathcal{D},\mathcal{A})}\left[\mathsf{CM}(x,p)\right]$, where $(x,p)\sim(\mathcal{D},\mathcal{A})$ denotes sampling a sequence $x$ and predictions $p$ from the joint distribution induced by $\mathcal{D}$ and $\mathcal{A}$.
@@
$\mathsf{SSCE}(x,p)\coloneq\operatorname*{\mathbb{E}}_{S\sim\mathsf{Unif}(2^{[T]})}\left[\mathsf{smCE}(x|_{S},p|_{S})\right]=\operatorname*{\mathbb{E}}_{y\sim\mathsf{Unif}(\{0,1\}^{T})}\left[\sup_{f\in\mathcal{F}}\sum_{t=1}^{T}y_{t}\cdot f(p_{t})\cdot(x_{t}-p_{t})\right].$
@@
With respect to distribution $\mathcal{D}\in\Delta(\{0,1\}^{T})$, the truthful forecaster is defined as $\mathcal{A}^{\mathrm{truthful}}(\mathcal{D})(b_{1},b_{2},\ldots,b_{t-1})\coloneq\Pr_{x\sim\mathcal{D}}\left[x_{t}=1\ \Big{|}\ x_{1:(t-1)}=b_{1:(t-1)}\right]$.
@@
The optimal error on distribution $\mathcal{D}\in\Delta(\{0,1\}^{T})$ with respect to calibration measure $\mathsf{CM}$ is defined as $\mathsf{OPT}_{\mathsf{CM}}(\mathcal{D})\coloneq\inf_{\mathcal{A}}\mathsf{err}_{\mathsf{CM}}(\mathcal{D},\mathcal{A})$, where $\mathcal{A}$ ranges over all deterministic forecasters.
@@ 3 Technical Overview
As we elaborate in Section 5, the crux of Step (2) is to control the expected deviation of a martingale $(M_{t})_{0\leq t\leq T}$ with respect to filtration $(\mathbb{F}_{t})$ by the its realized variance $\mathrm{Var}_{t}\coloneq\sum_{s=1}^{t}\operatorname*{\mathrm{Var}}\left[M_{s}|\mathbb{F}_{s-1}\right]$, which is highly non-trivial as the two processes $(M_{t})$ and $(\mathrm{Var}_{t})$ are correlated. In more detail, the filtration $(\mathbb{F}_{t})$ corresponds to the randomness in $x\sim\mathcal{D}$, while $(M_{t})$ tracks the biases in the predictions (on a subset of the time horizon) tested by a Lipschitz function. We note that such a bound would easily follow from “off-the-shelf” concentration inequalities for martingales (e.g., Freedman’s inequality [Fre75]), if the total realized variance $\mathrm{Var}_{T}$ were uniformly bounded. However, in general, $\mathrm{Var}_{T}$ may vary drastically, and directly applying these concentration inequalities would introduce an extra super-constant factor. Our workaround is a “doubling trick” that divides the time horizon into epochs, the realized variances in which grow exponentially. We then apply Freedman’s inequality to each epoch separately. In Section 5, we formulate a toy random walk problem that highlights this challenge and demonstrates our solution to it, which is of independent interest.
Similarly, as we show in Section 6, the crux of Step (3) is to establish another martingale inequality. We first show that for fixed $x$ and $p$, we have $\mathsf{SSCE}(x,p)=\Omega(\sqrt{N_{T}})$, where $N_{t}\coloneq\sum_{s=1}^{t}\mathbbm{1}\left[|x_{s}-p_{s}|\geq 1/2\right]$. Furthermore, over the randomness in $x\sim\mathcal{D}$, the realized variance process $(\mathrm{Var}_{t})$ defined above is shown to lower bound $(N_{t})$, i.e., $(N_{t}-\mathrm{Var}_{t})$ is a sub-martingale. However, the desired result requires the lower bound $\operatorname*{\mathbb{E}}\left[\sqrt{N_{T}}\right]\geq\Omega(1)\cdot\operatorname*{\mathbb{E}}\left[\sqrt{\mathrm{Var}_{T}}\right]$, which does not follow from $\operatorname*{\mathbb{E}}\left[N_{T}-\mathrm{Var}_{T}\right]\geq 0$ in general. This challenge necessitates a more careful analysis tailored to the specific properties of the processes $(N_{t})$ and $(\mathrm{Var}_{t})$.
@@ 4 Warmup: The Product Distribution Case
For distribution $\mathcal{D}=\prod_{t=1}^{T}\mathsf{Bernoulli}(p^{\star}_{t})$, take $\sigma^{2}\coloneq\operatorname*{\mathrm{Var}}_{x\sim\mathcal{D}}\left[\sum_{t=1}^{T}x_{t}\right]=\sum_{t=1}^{T}p^{\star}_{t}(1-p^{\star}_{t})$ as a complexity measure of the distribution of outcomes. We will show that $\mathsf{err}_{\mathsf{SSCE}}(\mathcal{D},\mathcal{A}^{\mathrm{truthful}}(\mathcal{D}))=O(\sigma+\log T)$ and $\mathsf{OPT}_{\mathsf{SSCE}}(\mathcal{D})=\Omega(\sigma)-O(1)$.
@@ 4.1 Upper Bound the SSCE of the Truthful Forecaster
where $\mathcal{G}_{\delta}\coloneq\{f_{\delta}-f_{\delta/2}:f_{\delta}\in\mathcal{F}_{\delta},f_{\delta/2}\in\mathcal{F}_{\delta/2},\|f_{\delta}-f_{\delta/2}\|_{\infty}\leq 3\delta/2\}$.
@@ 5 Upper Bound the SSCE of the Truthful Forecaster
To extend the proof strategy sketched in Section 4 to non-product distributions, the first challenge is to define an appropriate complexity measure of a general distribution $\mathcal{D}$. Consider the stochastic process $(\mathrm{Var}_{t})_{0\leq t\leq T}$ defined as $\mathrm{Var}_{t}\coloneq\sum_{s=1}^{t}p^{\star}_{s}(1-p^{\star}_{s})$, where
$p^{\star}_{t}\coloneq\operatorname*{\mathbb{E}}_{x^{\prime}\sim\mathcal{D}}\left[x^{\prime}_{t}\Big{|}x^{\prime}_{1:(t-1)}=x_{1:(t-1)}\right]$
@@
To control the expectation of each $W_{k}$, we note that the set $\mathcal{G}_{k}\coloneq\left\{\pi_{2^{-k}}(f)-\pi_{2^{1-k}}(f):f\in\mathcal{F}\right\}$ is of size at most $|\mathcal{F}_{2^{-k}}|\cdot|\mathcal{F}_{2^{1-k}}|$. Furthermore, every function $g\in\mathcal{G}_{k}$ satisfies
@@
Given a function $f:[0,1]\to[-1,1]$ and $y\in\{0,1\}^{T}$, consider the martingale $M_{t}(f,y)\coloneq\sum_{s=1}^{t}y_{s}\cdot f(p^{\star}_{s})\cdot(x_{s}-p^{\star}_{s})$ where $x\sim\mathcal{D}$. Then, for any finite family $\mathcal{G}$ of functions from $[0,1]$ to $[-1,1]$ and any $y\in\{0,1\}^{T}$, we have
@@
$\mathrm{Var}_{t}(\mathcal{I})\coloneq\sum_{s=1}^{t}p^{\star}_{s}(1-p^{\star}_{s})\cdot\mathbbm{1}\left[p^{\star}_{s}\in\mathcal{I}\right],$
where $x\sim\mathcal{D}$ and $p^{\star}_{t}\coloneq\Pr_{x^{\prime}\sim\mathcal{D}}\left[x^{\prime}_{t}=1|x^{\prime}_{1:(t-1)}=x_{1:(t-1)}\right]$. We define the epochs with respect to $\mathcal{I}$ as the sequence $\tau_{0},\tau_{1},\dots\in\mathbb{N}$ where $\tau_{0}=0$ and, for each $k\in[\left\lceil\log_{2}(T)\right\rceil+2]$,
@@ 5.2 Towards a Stronger Upper Bound
Given a function $f:[0,1]\to[-1,1]$, $y\in\{0,1\}^{T}$, and set $\mathcal{I}\subseteq[0,1]$, consider the martingale $M_{t}(f,y,\mathcal{I})\coloneq\sum_{s=1}^{t}y_{t}\cdot f(p^{\star}_{s})\cdot(x_{s}-p^{\star}_{s})\cdot\mathbbm{1}\left[p^{\star}_{s}\in\mathcal{I}\right]$, where $x\sim\mathcal{D}$, and $p_{t}^{\star}=\Pr_{x^{\prime}\sim\mathcal{D}}\left[x^{\prime}_{t}=1|x^{\prime}_{1:(t-1)}=x_{1:(t-1)}\right]$. Then, for any finite family $\mathcal{G}$ of functions from $[0,1]$ to $[-1,1]$, any $y\in\{0,1\}^{T}$, and any $\mathcal{I}\subseteq[0,1]$, we have